In a running event, a sprinter does 4.8 x 105 J of work and her internal energy decreases by 7.5 x 105 J. Determine the heat transferred between her body and surroundings during this event.
Answer #1

dq = du + dw

dq = -7.5 * 105 + 4.8 * 105

du = -ve sign show decrease internal energy

So,

dQ = -2.7 * 105J

So , amount of heat transfer between body and surroundings

Is dQ = 2.7 * 105J

plz upvote